OpenStudy (anonymous):

(3x^7+6x^4-5x+7)-(6x^4-5x^3+7) find the difference someone please help lol

jimthompson5910 (jim_thompson5910):

(3x^7+6x^4-5x+7)-(6x^4-5x^3+7) 3x^7+6x^4-5x+7-6x^4+5x^3-7 ... distribute now combine like terms

OpenStudy (anonymous):

wait what

jimthompson5910 (jim_thompson5910):

for example, 6x^4 and -6x^4 are like terms because they have the same exponent over the x

OpenStudy (anonymous):

so whats the answer

jimthompson5910 (jim_thompson5910):

so combining them gives you 6x^4 + (-6x^4) = 6x^4 - 6x^4 = 0x^4 = 0 that's just one pair of like terms

jimthompson5910 (jim_thompson5910):

see how I'm getting that?